### v1

@@ III-F Masked EEG Reconstruction (MER)
The reconstruction loss is calculated exclusively on masked patches to prevent trivial identity mapping. Let $\mathcal{M}=\{(c,s)\mid\mathbf{M}_{c,s}=1\}$ represent the set of masked indices, with $\mathbf{M}\in\{0,1\}^{C\times S}$ as the binary mask tensor. We utilize the Smooth L1 Loss (Huber loss) [huber1992robust]:
@@ III-G Neurodynamics Statistics Prediction (NSP)
While Masked EEG Reconstruction (MER) ensures local signal fidelity, it does not explicitly enforce the modeling of multi-scale dynamical structures defining brain states. To bridge this gap, we introduce Neurodynamics Statistics Prediction (NSP), a self-supervision paradigm grounded in dynamical systems theory [strogatz2024nonlinear, breakspear2017dynamic]. We conceptualize the brain as a high-dimensional nonlinear system whose macroscopic behavior is governed by a low-dimensional set of order parameters, representing stable, emergent collective variables [buzsaki2010neural].
@@
with observation function $\mathcal{H}(\cdot)$ and noise $\boldsymbol{\epsilon}(t)$. While $\mathbf{z}(t)$ is high-dimensional, its emergent dynamics are characterized by macroscopic order parameters $\mathbf{y}(t)=\Phi(\mathbf{z}(t))$, where $\Phi$ extracts stable features robust to microscopic perturbations [buzsaki2010neural].
@@
1. Relative Spectral Power: Local Oscillatory States. Spectral power distribution reflects the attractor landscape, with specific frequency bands linked to cognitive states [buzsaki2006rhythms]. We compute the Power Spectral Density (PSD) via real FFT:
@@
2. Phase-Locking Value (PLV): Functional Integration. Phase synchronization measures large-scale functional integration [varela2001brainweb]. For each band, we extract the analytic signal phase $\phi_{c}(t)$ via the Hilbert transform. The PLV between channels $i$ and $j$ is:
@@
3. Cross-Frequency Coupling (CFC): Hierarchical Coordination. CFC encodes coordination across temporal scales, such as theta–gamma interactions [canolty2006high]. For canonical pairs (e.g., theta–gamma), we extract the low-frequency phase $\phi_{\text{low}}(t)$ and high-frequency amplitude $A_{\text{high}}(t)$, computing the modulation strength:
@@
4. Sample Entropy: Dynamic Complexity. Signal complexity reflects the richness of neural dynamics [richman2000physiological]. Sample entropy quantifies regularity:
@@
By targeting these universal principles, NSP encourages the learning of representations invariant to subject-specific conditions [arjovsky2019invariant]. The total pretraining objective combines NSP with MER:
@@
TUEG (clean subset) [obeid2016temple] 256Hz 19 30s 1,109,545 9,246.2h
PhysioNet 2018 [Goldberger2000PhysioBankPA, ghassemi2018you] 200Hz 6 30s 872,089 7,267.4h
Raw EEG Data [raweegdata] 256Hz 64 5s 30,978 42.7h
Siena Scalp EEG Database [Goldberger2000PhysioBankPA, detti2020eeg] 256Hz 27 10s 50,749 140.8h
ds006171 [melcon2024perception, ds006171] 1024Hz 128 5s 29,397 40.8h
ds006317 [zhang2024chisco, ds006317] 1000Hz 109 5s 44,708 62.1h
ds006367 [ds006367] 1000Hz 28 8s 29,929 66.5h
ds006370 [ds006370] 1000Hz 28 5s 67,953 94.4h
ds006437 [ds006437] 256Hz 64 5s 12,069 16.8h
ds006446 [nashiro2024heart, kim2025electroencephalography, ds006446] 2048Hz 64 5s 12,970 18.0h
ds006466 [nashiro2024heart, kim2025electroencephalography, ds006466] 1000Hz 64 5s 94,392 131.1h
ds006480 [shou2020whole, ds006480] 1000Hz 64 5s 51,572 71.6h
ds006525 [ds006525] 250Hz 128 5s 4,080 5.7h
ds006547 [ds006547] 500Hz 63 5s 28,222 39.2h
@@
Emotion Recognition FACED [chen2023large] 250Hz 32 10s 10,332 9-class Sub 1-80 Sub 81-100 Sub 101-123
SEED-V [liu2021comparing] 1000Hz 62 1s 117,744 5-class Trials per session split 5:5:5
SEED-VII [jiang2024seed] 1000Hz 62 5s 55,561 7-class Trials per session split 10:5:5
Motor Imagery Classification PhysioNet-MI [goldberger2000physiobank, schalk2004bci2000] 160Hz 64 4s 9,837 4-class Sub 1-70 Sub 71-89 Sub 90-109
BCIC-IV-2a [brunner2008bci] 250Hz 22 4s 5,088 4-class Sub 1-5 Sub 6-7 Sub 8-9
SHU-MI [ma2022large] 250Hz 32 4s 11,988 2-class Sub 1-15 Sub 16-20 Sub 21-25
Sleep Staging ISRUC [khalighi2016isruc] 200Hz 6 30s 89,240 5-class Sub 1-80 Sub 81-90 Sub 91-100
Seizure Detection CHB-MIT [goldberger2000physiobank, shoeb2009application] 256Hz 16 10s 326,993 2-class Sub 1-19 Sub 20-21 Sub 22-23
Imagined Speech Classification BCIC2020-3 [jeong20222020] 256Hz 64 3s 6,000 5-class Predefined training, validation, and test
Mental Disorder Diagnosis MODMA [cai2022multi] 250Hz 128 15s 1,066 2-class 15MDD+18NC 4MDD+5NC 5MDD+6NC.
Vigilance Estimation SEED-VIG [min2017driver] 200Hz 17 8s 20,355 regression Sub 1-15 Sub 16-19 Sub 20-23
Workload Estimation MentalArithmetic [goldberger2000physiobank, zyma2019electroencephalograms] 500Hz 20 5s 1,707 2-class Sub 1-28 Sub 29-32 Sub 33-36
@@ IV-A1 Pretraining Datasets and Preprocessing
  • TUEG [obeid2016temple]: A large clinical EEG archive. Due to pervasive artifacts and unmarked noise in the full release, we curated a clean subset of high-quality 30-second segments consistent with CBraMod [wang2025cbramod].
  • PhysioNet 2018 [Goldberger2000PhysioBankPA, ghassemi2018you]: Sleep EEG from 994 subjects (training set only), annotated for sleep staging.
  • Raw EEG Data [raweegdata]: Recordings from cognitive task paradigms (information integration and rule-based categorization).
  • Siena Scalp EEG Database [Goldberger2000PhysioBankPA, detti2020eeg]: EEG from epilepsy patients, originally recorded with 31 channels but processed here using the standard 10–20 subset of 27 channels.
  • Ten OpenNeuro datasets (ds006171, ds006317, ds006367, ds006370, ds006437, ds006446, ds006466, ds006480, ds006525, ds006547) [openneuro]: A collection of contemporary cognitive and clinical EEG studies from the OpenNeuro platform, covering perceptual decision-making, emotional processing, resting-state, and motor imagery paradigms. These datasets vary widely in channel count (28–128) and sampling rate (250–2048 Hz), reflecting modern high-density and multi-center recording practices.
All recordings are preprocessed uniformly: (i) resampled to 200 Hz (except PhysioNet 2018, which is already at 200 Hz); (ii) band-pass filtered (0.3–75 Hz) to remove slow drifts and high-frequency noise; (iii) notch-filtered at 50/60 Hz to suppress power-line interference; (iv) segmented into non-overlapping fixed-length samples as specified per dataset; and (v) normalized to a scale of 100 µV (i.e., divided by 100), so that values predominantly lie in [1,1][-1,1], following the convention in LaBraM [jiang2024large].
@@ IV-A2 Pre-training Settings
DeeperBrain is implemented in Python 3.13.5 using PyTorch 2.8.0 with CUDA 12.9 for GPU acceleration. The model is pretrained on a single machine equipped with an Intel Xeon Gold 6226R CPU and one NVIDIA RTX A5000 GPU, completing training in approximately 7 hours. The input EEG signal is partitioned into non-overlapping spatiotemporal patches, each corresponding to a 1-second segment (200 time steps at 200 Hz). The backbone architecture consists of a 12-layer Transformer encoder with the following specifications: embedding dimension $D=200$, feed-forward hidden dimension $D_{\text{ff}}=800$, and 8 heads of self-attention. During pretraining, we adopt a random patch-wise masking strategy with a mask ratio of 50%, where masked patches are replaced by zero vectors in the raw voltage space. The model is trained for 2 epochs with a batch size of 16. We set the coefficients of two pretraining objectives as $\lambda_{\text{MER}}=\lambda_{\text{NSP}}=1.0$. Optimization is performed using the AdamW optimizer with a base learning rate of $5\times 10^{-4}$, weight decay of $5\times 10^{-2}$, and default $\beta_{1}=0.9$, $\beta_{2}=0.999$. No learning rate warmup is used, following the pre-normalization design’s stability [wang2019learning]. Gradient clipping with a maximum norm of 1.0 is applied to stabilize training. The code and pretrained model will be publicly available.
@@ IV-B1 Downstream Tasks and Datasets
To comprehensively evaluate the generalization capability of DeeperBrain, we selected 10 diverse downstream BCI tasks. Table II summarizes these tasks along with their corresponding datasets. Consistent with the pretraining configuration, all downstream EEG signals were resampled to 200 Hz, and the patch duration was fixed at 1 second (200 data points). To ensure fair comparison, we adopted standard data splitting protocols from LaBraM [jiang2024large] or CBraMod [wang2025cbramod] whenever available.
@@ IV-B3 Compared Methods
We compare DeeperBrain with both non-foundation-model and foundation-model baselines on all the downstream BCI tasks. We adopt EEGNet [lawhern2018eegnet] and EEGConformer [song2022eeg] as non-foundation-model baselines. Concurrently, we use four very strong existing methods, LaBraM [jiang2024large], CBraMod [wang2025cbramod], CSBrain [zhou2025csbrain] and REVE [ouahidi2025reve] as the foundation-model baselines.
@@ VI Conclusion
VII Biography Section
[Uncaptioned image] Jiquan Wang received the Ph.D. degree in Computer Science and Technology from Zhejiang University, Hangzhou, China, in 2025. He is currently a research fellow at the State Key Laboratory of Brain-Machine Intelligence, Zhejiang University. His research interests include EEG decoding, brain-computer interfaces and artificial intelligence. He received the 2025 ACM Hangzhou Outstanding Doctoral Dissertation Award.
[Uncaptioned image] Sha Zhao received the Ph.D. degree from Zhejiang University, Hangzhou, China, in 2017. She is currently a Postdoctoral Research Fellow with the College of Computer Science and Technology, Zhejiang University. She visited the Human-Computer Interaction Institute, Carnegie Mellon University, Pittsburgh, PA, USA, as a visiting Ph.D. student from 2015 to 2016. Her research interests include brain-machine interfaces, data mining and machine learning. Dr. Zhao received the Best Paper Award of ACM UbiComp’16.
[Uncaptioned image] Yangxuan Zhou received the B.S. degree in Electronic Information from Shanghai University, Shanghai, China, in 2022. He is currently working toward the Ph.D. degree of Computer Science and Technology with Zhejiang University, Hangzhou, China. His research interests include EEG decoding, Brain-Computer Interfaces, Deep Learning, and Artificial Intelligence.
[Uncaptioned image] Yiming Kang received the B.S. degree in Artificial Intelligence from Soochow University, Suzhou, China, in 2024. He has been pursuing the M.S. degree in Electronic Information with Zhejiang University, Hangzhou, China, since 2025. His research primarily focuses on multimodal brain foundation model in the field of non-invasive brain-computer interfaces (BCI).
[Uncaptioned image] Shijian Li received the Ph.D. degree from Zhejiang University, Hangzhou, China, in 2006. In 2010, he was a Visiting Scholar with the Institute Telecom SudParis, Évry, France. He is currently with the College of Computer Science and Technology, Zhejiang University. He has published over 40 papers. His research interests include sensor networks, ubiquitous computing, and social computing. Dr. Li serves as an Editor for the International Journal of Distributed Sensor Networks and as a reviewer or the PC member of over ten conferences.
[Uncaptioned image] Gang Pan (Senior Member, IEEE) received the B.Eng. and Ph.D. degrees from Zhejiang University, Hangzhou, China, in 1998 and 2004, respectively. He is currently a Professor with the Department of Computer Science, and the Deputy Director of the State Key Lab of CAD&CG, Zhejiang University. From 2007 to 2008, he was a Visiting Scholar with the University of California, Los Angeles, CA, USA. His current interests include brain-inspired computing, brain-machine interfaces, artificial intelligence and pervasive computing.

### v2

@@ III-F Masked EEG Reconstruction (MER)
The reconstruction loss is calculated exclusively on masked patches to prevent trivial identity mapping. Let $\mathcal{M}=\{(c,s)\mid\mathbf{M}_{c,s}=1\}$ represent the set of masked indices, with $\mathbf{M}\in\{0,1\}^{C\times S}$ as the binary mask tensor. We utilize the Smooth L1 Loss (Huber loss) [Huber, 1992]:
@@ III-G Neurodynamics Statistics Prediction (NSP)
While Masked EEG Reconstruction (MER) ensures local signal fidelity, it does not explicitly enforce the modeling of multi-scale dynamical structures defining brain states. To bridge this gap, we introduce Neurodynamics Statistics Prediction (NSP), a self-supervision paradigm grounded in dynamical systems theory [Strogatz, 2024; Breakspear, 2017]. We conceptualize the brain as a high-dimensional nonlinear system whose macroscopic behavior is governed by a low-dimensional set of order parameters, representing stable, emergent collective variables [Buzsáki, 2010].
@@
with observation function $\mathcal{H}(\cdot)$ and noise $\boldsymbol{\epsilon}(t)$. While $\mathbf{z}(t)$ is high-dimensional, its emergent dynamics are characterized by macroscopic order parameters $\mathbf{y}(t)=\Phi(\mathbf{z}(t))$, where $\Phi$ extracts stable features robust to microscopic perturbations [Buzsáki, 2010].
@@
1. Relative Spectral Power: Local Oscillatory States. Spectral power distribution reflects the attractor landscape, with specific frequency bands linked to cognitive states [Buzsáki, 2006]. We compute the Power Spectral Density (PSD) via real FFT:
@@
2. Phase-Locking Value (PLV): Functional Integration. Phase synchronization measures large-scale functional integration [Varela et al., 2001]. For each band, we extract the analytic signal phase $\phi_{c}(t)$ via the Hilbert transform. The PLV between channels $i$ and $j$ is:
@@
3. Cross-Frequency Coupling (CFC): Hierarchical Coordination. CFC encodes coordination across temporal scales, such as theta–gamma interactions [Canolty et al., 2006]. For canonical pairs (e.g., theta–gamma), we extract the low-frequency phase $\phi_{\text{low}}(t)$ and high-frequency amplitude $A_{\text{high}}(t)$, computing the modulation strength:
@@
4. Sample Entropy: Dynamic Complexity. Signal complexity reflects the richness of neural dynamics [Richman and Moorman, 2000]. Sample entropy quantifies regularity:
@@
By targeting these universal principles, NSP encourages the learning of representations invariant to subject-specific conditions [Arjovsky et al., 2019]. The total pretraining objective combines NSP with MER:
@@
TUEG (clean subset) [Obeid and Picone, 2016] 256Hz 19 30s 1,109,545 9,246.2h
PhysioNet 2018 [Goldberger et al., 2000b; Ghassemi et al., 2018] 200Hz 6 30s 872,089 7,267.4h
Raw EEG Data [Trujillo, 2020] 256Hz 64 5s 30,978 42.7h
Siena Scalp EEG Database [Goldberger et al., 2000b; Detti et al., 2020] 256Hz 27 10s 50,749 140.8h
ds006171 [Melcón et al., 2024, 2025] 1024Hz 128 5s 29,397 40.8h
ds006317 [Zhang et al., 2024b, 2025] 1000Hz 109 5s 44,708 62.1h
ds006367 [xx, 2025] 1000Hz 28 8s 29,929 66.5h
ds006370 [X, 2025] 1000Hz 28 5s 67,953 94.4h
ds006437 [anonymous, 2025] 256Hz 64 5s 12,069 16.8h
ds006446 [Nashiro et al., 2024; Kim et al., 2025b; Kinley and Becker, 2025] 2048Hz 64 5s 12,970 18.0h
ds006466 [Nashiro et al., 2024; Kim et al., 2025b, c] 1000Hz 64 5s 94,392 131.1h
ds006480 [Shou et al., 2020; Kim et al., 2025a] 1000Hz 64 5s 51,572 71.6h
ds006525 [Neuroimaging and ar the University of Oklahoma, 2025] 250Hz 128 5s 4,080 5.7h
ds006547 [Ghaffari et al., 2025] 500Hz 63 5s 28,222 39.2h
@@
Emotion Recognition FACED [Chen et al., 2023] 250Hz 32 10s 10,332 9-class Sub 1-80 Sub 81-100 Sub 101-123
SEED-V [Liu et al., 2021b] 1000Hz 62 1s 117,744 5-class Trials per session split 5:5:5
SEED-VII [Jiang et al., 2024a] 1000Hz 62 5s 55,561 7-class Trials per session split 10:5:5
Motor Imagery Classification PhysioNet-MI [Goldberger et al., 2000a; Schalk et al., 2004] 160Hz 64 4s 9,837 4-class Sub 1-70 Sub 71-89 Sub 90-109
BCIC-IV-2a [Brunner et al., 2008] 250Hz 22 4s 5,088 4-class Sub 1-5 Sub 6-7 Sub 8-9
SHU-MI [Ma et al., 2022] 250Hz 32 4s 11,988 2-class Sub 1-15 Sub 16-20 Sub 21-25
Sleep Staging ISRUC [Khalighi et al., 2016] 200Hz 6 30s 89,240 5-class Sub 1-80 Sub 81-90 Sub 91-100
Seizure Detection CHB-MIT [Goldberger et al., 2000a; Shoeb, 2009] 256Hz 16 10s 326,993 2-class Sub 1-19 Sub 20-21 Sub 22-23
Imagined Speech Classification BCIC2020-3 [Jeong et al., 2022] 256Hz 64 3s 6,000 5-class Predefined training, validation, and test
Mental Disorder Diagnosis MODMA [Cai et al., 2022] 250Hz 128 15s 1,066 2-class 15MDD+18NC 4MDD+5NC 5MDD+6NC.
Vigilance Estimation SEED-VIG [Min et al., 2017] 200Hz 17 8s 20,355 regression Sub 1-15 Sub 16-19 Sub 20-23
Workload Estimation MentalArithmetic [Goldberger et al., 2000a; Zyma et al., 2019] 500Hz 20 5s 1,707 2-class Sub 1-28 Sub 29-32 Sub 33-36
@@ IV-A1 Pretraining Datasets and Preprocessing
  • TUEG [Obeid and Picone, 2016]: A large clinical EEG archive. Due to pervasive artifacts and unmarked noise in the full release, we curated a clean subset of high-quality 30-second segments consistent with CBraMod [Wang et al., 2025a].
  • PhysioNet 2018 [Goldberger et al., 2000b; Ghassemi et al., 2018]: Sleep EEG from 994 subjects (training set only), annotated for sleep staging.
  • Raw EEG Data [Trujillo, 2020]: Recordings from cognitive task paradigms (information integration and rule-based categorization).
  • Siena Scalp EEG Database [Goldberger et al., 2000b; Detti et al., 2020]: EEG from epilepsy patients, originally recorded with 31 channels but processed here using the standard 10–20 subset of 27 channels.
  • Ten OpenNeuro datasets (ds006171, ds006317, ds006367, ds006370, ds006437, ds006446, ds006466, ds006480, ds006525, ds006547) [OpenNeuro, 2025]: A collection of contemporary cognitive and clinical EEG studies from the OpenNeuro platform, covering perceptual decision-making, emotional processing, resting-state, and motor imagery paradigms. These datasets vary widely in channel count (28–128) and sampling rate (250–2048 Hz), reflecting modern high-density and multi-center recording practices.
All recordings are preprocessed uniformly: (i) resampled to 200 Hz (except PhysioNet 2018, which is already at 200 Hz); (ii) band-pass filtered (0.3–75 Hz) to remove slow drifts and high-frequency noise; (iii) notch-filtered at 50/60 Hz to suppress power-line interference; (iv) segmented into non-overlapping fixed-length samples as specified per dataset; and (v) normalized to a scale of 100 µV (i.e., divided by 100), so that values predominantly lie in [1,1][-1,1], following the convention in LaBraM [Jiang et al., 2024b].
@@ IV-A2 Pre-training Settings
DeeperBrain is implemented in Python 3.13.5 using PyTorch 2.8.0 with CUDA 12.9 for GPU acceleration. The model is pretrained on a single machine equipped with an Intel Xeon Gold 6226R CPU and one NVIDIA RTX A5000 GPU, completing training in approximately 7 hours. The input EEG signal is partitioned into non-overlapping spatiotemporal patches, each corresponding to a 1-second segment (200 time steps at 200 Hz). The backbone architecture consists of a 12-layer Transformer encoder with the following specifications: embedding dimension $D=200$, feed-forward hidden dimension $D_{\text{ff}}=800$, and 8 heads of self-attention. During pretraining, we adopt a random patch-wise masking strategy with a mask ratio of 50%, where masked patches are replaced by zero vectors in the raw voltage space. The model is trained for 2 epochs with a batch size of 16. We set the coefficients of two pretraining objectives as $\lambda_{\text{MER}}=\lambda_{\text{NSP}}=1.0$. Optimization is performed using the AdamW optimizer with a base learning rate of $5\times 10^{-4}$, weight decay of $5\times 10^{-2}$, and default $\beta_{1}=0.9$, $\beta_{2}=0.999$. No learning rate warmup is used, following the pre-normalization design’s stability [Wang et al., 2019]. Gradient clipping with a maximum norm of 1.0 is applied to stabilize training. The code and pretrained model will be publicly available.
@@ IV-B1 Downstream Tasks and Datasets
To comprehensively evaluate the generalization capability of DeeperBrain, we selected 10 diverse downstream BCI tasks. Table II summarizes these tasks along with their corresponding datasets. Consistent with the pretraining configuration, all downstream EEG signals were resampled to 200 Hz, and the patch duration was fixed at 1 second (200 data points). To ensure fair comparison, we adopted standard data splitting protocols from LaBraM [Jiang et al., 2024b] or CBraMod [Wang et al., 2025a] whenever available.
@@ IV-B3 Compared Methods
We compare DeeperBrain with both non-foundation-model and foundation-model baselines on all the downstream BCI tasks. We adopt EEGNet [Lawhern et al., 2018] and EEGConformer [Song et al., 2022] as non-foundation-model baselines. Concurrently, we use four very strong existing methods, LaBraM [Jiang et al., 2024b], CBraMod [Wang et al., 2025a], CSBrain [Zhou et al., 2025b] and REVE [Ouahidi et al., 2025] as the foundation-model baselines.